### v2

# Prompt Perturbations Reveal Human-Like Biases in LLM Survey Responses

Abstract
Large Language Models (LLMs) are increasingly used as proxies for human subjects in social science surveys, but their reliability and susceptibility to known response biases are poorly understood. This paper investigates the response robustness of LLMs in normative survey contexts—we test nine diverse LLMs on questions from the World Values Survey (WVS), applying a comprehensive set of 11 perturbations to both question phrasing and answer option structure, resulting in over 167,000 simulated interviews. In doing so, we not only reveal LLMs’ vulnerabilities to perturbations but also show that all tested models exhibit a consistent, but variably intense, recency bias, disproportionately favoring the last-presented answer option. While larger models are generally more robust, all models remain sensitive to semantic variations like paraphrasing and to combined perturbations. By applying a set of perturbations, we reveal that LLMs partially align with survey response biases identified in humans. This underscores the critical importance of prompt design and robustness testing when using LLMs to generate synthetic survey data.
Prompt Perturbations Reveal Human-Like Biases
in LLM Survey Responses
Jens Rupprecht1, Georg Ahnert1, and Markus Strohmaier1,2,3 1University of Mannheim, Mannheim 2GESIS - Leibniz Institute for the Social Sciences, Cologne 3Complexity Science Hub, Vienna
1 Introduction
Problem
Large Language Models (LLMs) are increasingly being used as proxies for human subjects in social science research, particularly for generating synthetic responses to survey questions (Argyle et al., 2023; Bisbee et al., 2024, inter alia). This application holds promise for augmenting or replacing costly human data collection, but the reliability of these synthetic respondents and to what extent they overlap with human responses remain open questions. In particular, survey methodology research has found that human responses are sensitive to subtle variations in question and answer phrasing that lead to well-known response biases Krosnick (1991) and it remains unclear whether LLMs, trained on vast amounts of human text, exhibit the same vulnerabilities.
Approach
In this paper we present a large-scale empirical study where we investigate the response behavior and robustness of nine different LLMs to normative questions derived from the World Value Survey (WVS; Haerpfer et al., 2022). We systematically apply eleven perturbations, targeting both the question phrasing and the structure of the answer options, as shown in Figure 1. Our research questions are:
  1. 1.
    Do prompt perturbations negatively affect the robustness of LLMs when answering closed-ended, normative survey questions?
  2. 2.
    Do LLMs exhibit human-like response biases when answering closed-ended, normative survey questions?
We find a consistent recency bias across all tested models in different strengths, where the last-presented answer option is disproportionately favored. Larger models generally exhibit more stable response patterns, but even the largest models are sensitive to changes in question phrasing.
Contribution
We conduct an experimental evaluation with over 167,000 interviews across nine distinct LLMs, varying in size and origin. We develop and apply a comprehensive set of 11 perturbations, targeting classic survey biases (e.g., response order, scale structure) with common textual variations (e.g., typos, paraphrasing). This makes it a useful baseline to test newly developed LLMs on the same task to put the results into perspective. We provide a detailed analysis of LLM response stability, showing that while some models are more robust than others, all are susceptible to specific types of perturbations. This work underscores the importance of careful prompt and Q&A design when using LLMs as a resource for synthetic survey responses, and provides a framework for evaluating their robustness in survey contexts. [^1] The material can be accessed at https://github.com/dess-mannheim/Survey-Response-Biases-in-LLMs.git.
Figure 1: The Interview Process. The figure shows an example of a bias perturbation (e.g. reversed option order) and a non-bias perturbation (e.g. typos in the question). Each model is interviewed 25 times with each combination.
2 Related Work
Our work builds on two main streams of research: (1) survey methodology from the social sciences, which documents human response biases, and (2) recent studies in computer science on the robustness and biases of LLM’s synthetic survey response generation.
Human Survey Response Biases
Research in the social sciences has shown that how a survey question is asked can be as important as what is asked. Respondents often engage in "satisficing" rather than "optimizing", choosing a satisfactory answer with minimal cognitive effort instead of carefully formulating an optimal one Krosnick (1991). This can lead to systematic biases. For example, the order in which the answer options are presented can induce primacy (favoring early options in visual surveys) or recency (favoring later options in oral surveys) biases Krosnick and Alwin (1987). The presence or absence of a middle option or a "don’t know" category can trigger a central tendency bias or opinion floating, respectively Hollingworth (1910); Koch and Blohm (2016). In the first, if a central category is available on the answer scale, humans tend to choose the central category, whereas opinion floating indicates that responses are redistributed to central categories if a refusal category is missing Tjuatja et al. (2024). In addition, priming effects, where the preceding context influences subsequent responses, are a well documented phenomenon Bargh et al. (1996). We draw on these past findings to design perturbations testing whether LLMs exhibit similar human-like response patterns.
LLMs as Survey Respondents
Recent studies explored LLMs as substitutes for human survey participants to generate synthetic data. They found that LLMs can replicate average public opinion on political topics, but often with less variance than human samples (Argyle et al., 2023; Bisbee et al., 2024; Von Der Heyde et al., 2025, inter alia). Others have found that LLM responses can be sensitive to prompting, revealing cultural and demographic biases Geng et al. (2024).
Our work is related to that of Tjuatja et al. (2024), who were among the first to systematically explore human-like response biases in LLMs. They investigated acquiescence, response order, opinion floating, and scale structure effects. Our study extends their work by: (1) using a different, globally diverse survey (the World Values Survey); (2) testing a wider range of LLMs of varying sizes and developers; and (3) incorporating a broader set of perturbations on both answer and question phrasing, such as keyboard typos, paraphrasing, synonyms, priming as well as a combined interaction of two perturbations.
LLM Robustness to Perturbations
Other researchers have evaluated the general robustness of LLMs to noisy or varied inputs on different tasks. They have shown that even state-of-the-art models can be sensitive to minor changes in the prompt. These perturbations range from the character level, such as typos created by swapping, inserting, or replacing letters Moradi and Samwald (2021); Gan et al. (2024), to word- or sentence-level, such as replacing words with synonyms or paraphrasing entire sentences Qiang et al. (2024). A common finding is that character-level noise can significantly degrade performance, even in large models Gan et al. (2024). The combination of multiple perturbations can even have a more negative effect Dong et al. (2023). Although this research has primarily focused on knowledge-based or reasoning tasks, we adapt these perturbation techniques to the context of normative surveys to assess response stability where no single "correct" answer exists.
Evaluation and Prompting
Finally, our work is guided by research on identified ways for evaluating LLMs on multiple-choice tasks. Studies have shown that evaluation results can be highly sensitive to prompt format, e.g. if LLMs face an open- or closed-ended response, and forcing technique. However, forcing a model to choose from a predefined set of options is often necessary to obtain valid responses, as unconstrained answers can differ substantially Röttger et al. (2024). The returned response labels might differ significantly when a LLM has the option to generate text output before returning the response label due to their auto-regressive nature. Furthermore, relying on the model’s first predicted token can misrepresent its full textual output Wang et al. (2024).
3 Methods
First, we select a subset of 62 questions representing a sample of different thematic categories, each question in the category sharing the same answer options. These normative, value-oriented Q&A pairs are taken from the WVS’s core variables Haerpfer et al. (2022), excluding all sociodemographic variables. On each Q&A pair we perform all eleven perturbations.
Figure 1 illustrates two exemplary perturbations and the interview process. In total, we perform five perturbations on the answer options as well as five perturbations on the question phrasing of the chosen subset questionnaire. We further include one interaction of two perturbations - one on the answer option and one on the question.
Interviews with the original and each perturbed Q&A pair are carried out 25 times with nine different LLMs. In total, we conducted 167,400 interviews, 18,600 with each model. We compare the distributions of the responded labels for all perturbations through entropy and KL-divergence to the baseline response distribution to the original Q&A pairs. Primacy bias is further examined by comparing the response frequencies of the first and last answer option in the list, whereas opinion floating bias and central tendency bias are tested by checking the shift of responses toward or away from the center of the answer option scale.
3.1 Experimental Setup
Survey Data
The questions and answer options are sourced from the WVS Wave 7 (2017-2022), a comprehensive cross-national survey on human beliefs and values Haerpfer et al. (2022). From the 259 core variables concerning values and norms, we selected a subset of 62 questions. This subset represents 10 distinct thematic categories, including Trust in People, Confidence in Institutions, Moral Justifiability, and Perception of Democracy, ensuring a diverse range of topics and answer scale formats (e.g., 3-point to 10-point scales).
Models
To ensure our findings are not specific to a single model architecture or developer, we selected nine instruction-tuned LLMs, varying in size, developer, and origin. This selection aims to establish a degree of external validity for our results and includes proprietary and open-source LLMs. The following models were interviewed: Gemini-1.5-Pro, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Llama-3.2-1B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3.5-mini-instruct, Qwen2.5-7B-Instruct and Yi-1.5-6B-Chat.
Infrastructure
Experiments were carried out on a high-performance computing cluster (bwUniCluster 2.0) and a local server equipped with NVIDIA H100 (80GB) GPUs. The total runtime on bwUniCluster 2.0 for one model’s 18,600 interviews, e.g. Llama-3.1-8B-Instruct including all perturbed and original Q&As, was ca. 35 minutes with approximately 0.11 seconds per interview. To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct. Smaller models were run without quantization. Gemini-1.5-Pro was accessed via its official API.
3.2 Perturbation Design
We designed two categories of perturbations to test model robustness: (1) bias-inducing alterations to the answer options, based on survey methodology research and known to induce biased responses in humans, and (2) non-bias alterations to the question phrasing, mimicking common textual variations and errors. Appendix Table 1 provides examples as well as references for all the perturbations. For each of the 62 Q&A pairs, we created the following eleven perturbed versions.
Bias Perturbations
These perturbations manipulate the provided answer choices to test for known survey response biases identified in human subjects.
  • Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
  • Missing Refusal: The “Don’t know” or refusal option is removed from the list of choices.
  • Odd/Even Scale Transformation: For scales with an even number of options, we use Gemini-1.5-flash to generate a semantically appropriate middle category, transforming it into an odd-numbered scale (e.g., adding ‘Neutral‘). Conversely, for odd-numbered scales, we remove the middle category to create an even-numbered scale and adjust the integer label.
  • Emotional Priming: A sentence designed to elicit a response is appended to the prompt after the answer options: ‘This is very important to my research! You better do not refuse the answer.’
Non-Bias Perturbations
These perturbations modify the question text to assess robustness to stylistic variations and typos. Typically humans are unaffected by such subtle changes in the question phrasing and are still able to understand the question’s meaning.
  • Typographical Errors: We introduce three types of typos: Letter Swap (swapping two letters in a random word), Key Typo (replacing a character with a random one), and Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
  • Semantic Variations: We use Gemini-1.5-flash to create two semantic variations while preserving the original meaning: First, by Synonym Replacement where five words in the original question are replaced with synonyms. Second, through Paraphrasing the entire question is rephrased.
We manually validated all LLM-generated perturbations (paraphrases, synonyms, odd-scale options) on our 62-question subset to correct errors and ensure their semantic integrity.
Interaction Effect
To study combined effects, we created one additional condition that pairs a paraphrased question with reversed-order answer options.
3.3 Interview Procedure and Data Collection
Prompting
To ensure internal validity, we used a single, consistent prompt structure for all interviews. The prompt was designed to be clear and direct, instructing the model to select a single option and return only its label.
Answer the following question: {question}
This is a list of possible answer options:
{answer_options}
You must pick one of the answer options. Only answer with the label.
Taking into account the findings of Röttger et al. (2024) we use a combination of their forced choice prompts that rendered the most valid responses, which means that the model is correctly choosing one available option from the given answer options. In an unforced prompt setting, the model did not follow the instructions and only rarely returned valid response labels, making a statistical evaluation of the results infeasible.
Data Collection
Each of the 9 models was presented with 12 experimental conditions (1 original + 11 perturbations) for each of the 62 selected WVS questions. To obtain a stable distribution of responses and enable statistical analysis, we repeated each unique model-Q&A-perturbation combination 25 times. This resulted in a total of $9\times 62\times 12\times 25=167,400$ interviews (cf. Figure 1).
Response Extraction and Validation
Preliminary tests revealed that smaller LLMs often fail to perfectly follow the instruction to "answer only with the label", returning conversational filler or explanations alongside their choice. To ensure accurate data for analysis, we developed a robust extraction pipeline.
We compared two main approaches. First, Gemini-1.5-Pro, Llama-3.1-8B, and Qwen2.5-7B were prompted, and a regular expression was designed to extract the answer labels. Based on multiple conditions, e.g. if the given answer label is part of the original answer options or that only one response is provided, the methods should highlight which technique is the most promising in extracting valid responses and handling possible edge cases of model responses.
Figure 2: Combined share of invalid and refusal responses by model and perturbation. Larger models like Llama-70B, Phi-3.5, and Mistral are highly reliable. Qwen2.5 shows high non-response rates, especially for typo-based perturbations and sensitive topics.
We manually labeled these extraction methods on a random sample of responses for validation. The LLM-based methods achieved accuracies between 77% and 97.5%, with the largest model Gemini-1.5-Pro performing best. However, our refined regular expression achieved the overall best extraction success on the validation set as it correctly extracted all responses in our validation set. Consequently, we used this regular expression to process all remaining 167,400 model responses.
4 Results
This section presents the results of our experiments, focusing on three key areas: (1) the models’ adherence to interview instructions, (2) their general and overall robustness to various input perturbations, and (3) their susceptibility to human-like survey response biases revealed by interviews with prompt perturbations.
4.1 Interview Adherence and Refusal Rates
Overall, the models demonstrated high adherence to the prompt instructions, with an average of 96% of interviews yielding an extractable and valid answer that was part of the given answer options. However, performance varied significantly across models. Larger models such as Llama-3.3-70B and Gemini-1.5-Pro, but also Phi-3.5-mini and Mistral are very reliable response generators and followed the instructions well while returning little to no incorrect or no answer label. In contrast, other models, particularly smaller Llama models like Llama-3.2-3B (83.6%) and Qwen2.5-7B, were more likely to produce invalid responses that did not follow instructions.
We combined invalid responses with explicit refusals (i.e., choosing the Don’t know option) to measure overall non-response rates, as shown in Figure 2. Llama-3.3-70B, Phi-3.5, and Mistral-7B consistently provided on-scale answers, with non-response rates typically below 10%. Conversely, Qwen2.5-7B and Llama-3.1-8B exhibited high non-response rates, often exceeding 30%.
Notably, we observed topic-specific sensitivity. For questions regarding the Perception of Elections, Qwen2.5-7B failed to provide a valid, on-scale answer in 91.3% of cases, even for the original, unperturbed questions. This might suggest the presence of strong content-based guardrails or restrictions in certain models (cf. Figure 8).
Figure 3: Evidence of recency bias across all models. The bars show the frequency of choosing the same answer option (e.g., “Very important”) when it is presented first vs. last. All models are significantly more likely to select an option when it appears at the end of the list.
4.2 Consistency and Robustness to Question and Answer Perturbations
We distinguish between response consistency (the tendency of a model to give the same answer to the same prompt, measured by entropy) and response robustness (the tendency to maintain a similar answer distribution under perturbation, measured by KL-divergence). A KL divergence of zero indicates a perfect match and thus full robustness against the input perturbation, whereas a high entropy value indicates very inconsistent response behavior.
Effect of Model Size and Consistency
First, we found that model size is in an inverse relationship with response consistency; smaller models exhibited higher entropy and standard deviation when asked the same question multiple times, indicating more random response behavior.
When assessing robustness to perturbations, we found a clear relationship with model size: larger models tend to be more robust. Figure 4 shows the percentage of questions for which the models produced a perfectly identical response distribution (KL-Div = 0) despite perturbations. Llama-3.3-70B and Gemini-1.5-Pro were the most robust, often replicating their original answers in over 50% of cases. The smaller Llama models were the least robust, with Llama-3.2-1B perfectly replicating its answers in fewer than 5% of cases on average. This suggests that scale is a key factor in achieving stable response behavior in synthetic response generation.
Figure 4: Share of fully robust responses (KL-Divergence = 0) by model and perturbation type. Top: Non-bias perturbations. Bottom: Bias perturbations. A clear trend emerges: larger models (Llama-70B, Gemini) are substantially more robust than smaller models (Llama-1B, Llama-3B). Certain perturbations, like the combined interaction effect, paraphrasing, and reversing the scale, are challenging for all models.
Effect of Perturbation Type
Further, Figure 4 highlights the share of fully robust responses (Kullback-Leibler Divergence = 0) across all questions by perturbation and LLM. It shows that certain perturbations had a greater impact on robustness across all models.
  • Combined Perturbations: The interaction of two perturbations (paraphrased question + reversed answers) has the most bewildering effect on the responses, causing the lowest robustness scores for all models except Phi-3.5-mini.
  • Semantic vs. Lexical Changes: Paraphrasing the question reduced robustness more than replacing individual words with synonyms in most LLMs. These findings are in line with Moradi and Samwald (2021) who found that models trained on larger corpora are more robust when words are replaced by their synonyms.
  • Typographical Errors: Randomly replacing characters (Key Typo) or using adjacent keys (Keyboard Typo) was more robustness-harming than simply swapping two letters within a word (Letter Swap).
  • Answer Option Changes: Reversing the answer scale or changing it from odd to even (or vice versa) had a more negative impact on the robustness of responses than removing the refusal option or adding emotional priming.
Effect of Scale Length
We also observed that robustness is sensitive to the complexity of the task. For nearly all models, the share of fully robust responses decreased as the length of the answer scale, i.e. answer options, increased. For example, models were less likely to replicate their exact response distributions on a 10-point scale compared to a 4-point scale, indicating that a larger decision space can make LLMs more susceptible to perturbations. Figures 12 and 13 suggest that for most LLMs, except Gemini-1.5-pro, the size of the answer option scale has an impact on response robustness comparing the share of fully robust responses on e.g. the 4- and 10-point scale. This suggests that the larger the answer scale, the less likely models can reproduce the responses they gave in the original Q&A phrasing, under perturbed settings.
4.3 Evidence of Human-like Survey Biases
With some of the perturbations, we are able to go beyond LLMs robustness and consistency but also analyze whether LLMs exhibit human-like survey response biases. We find evidence for some human-like biases.
Recency Bias
Contrary to the initially hypothesized primacy bias, we found weak to strong, but consistent indications of a recency bias across all nine models. When we reversed the order of the answer scale, the probability to choose the first option plummeted, while the probability to choose the last option (which is the semantically identical first option in the original Q&A) increased strongly, ceteris paribus. As shown in Figure 3, this effect was substantial, with the choice frequency of the semantically same option increasing by over 2000% for Llama-3.1-8B when moved to the last position while all other configurations, such as question and prompt phrasing, were held constant. This indicates that LLMs, similar to human respondents in oral surveys, might overemphasize the final options they process.
Opinion Floating and Central Tendency
The effects of removing the refusal option (opinion floating) or providing an explicit middle category (central tendency) were highly model-dependent, often correlated with model size (Figures 9(a) and 9(b)). For opinion floating, larger models like Llama-70B, Gemini, but also Phi-3.5 were largely robust, showing minimal shifts in their response distributions. Smaller models, particularly Qwen and Llama-8B, showed a weak tendency to shift responses toward the scale’s center when the refusal option was absent. Here, we expect that models redistribute their original non-responses to the center of the answer scale to maintain their indecisiveness, which is also known as opinion floating bias. Similarly, for central tendency, larger models (Llama-70B, Gemini, Mistral) consistently shifted their mean response closer to the center when an explicit middle option was provided. Binomial tests underlined that the middle option was selected significantly more often than expected under a uniform distribution, especially on longer scales. Smaller models, however, showed inconsistent effects or were completely unaffected.
Emotional Priming
The impact of adding an emotional priming statement (“This is very important to my research!”) was also model-dependent. For larger models (Llama-70B, Gemini, Mistral), it either had no effect or slightly decreased the rate of refusal responses, suggesting they correctly interpreted the intent of the priming statement. Conversely, for the two Chinese models, Qwen2.5-7B and Yi-1.5-6B, the priming text even increased the share of refusal responses across most topics. No clear relationship can be drawn from these findings due to the inconsistent behavior across models.
5 Discussion and Conclusion
We found that perturbations can be an insightful approach to identify human-like survey response biases. For example, the same answer option is more likely chosen if it comes as the last mentioned option than if it was the first answer option. This consistent shift of response distribution to the last answer option across LLMs suggests a recency bias rather than a primacy bias. Although this is not valid across all inspected models, excluding an explicit refusal category, and especially adding a middle category or odd scale instead of an even scale, shifts the mean response more toward the original central point of the scale. Thus, a central tendency bias could only be identified for specific models across all scale types, whereas none of the LLMs consistently mirrors a opinion floating bias.
Furthermore, the variety of perturbation allows us to gain insights into the robustness of LLMs as some models are more sensitive and some perturbations are more robustness-harming than others. For instance, swapping letters within a word has less negative impact than introducing random or keyboard-adjacent characters. This might be explained by the fact that letter swaps are more likely when typing and therefore might potentially take a greater part of the training data. This possibly makes the LLM more resilient to this perturbation compared to exchanging characters with random others. Combining two perturbations has the strongest negative impact on robustness, whereas synonyms tend to be less confusing than paraphrasing.
Future work should also focus on the impact of persona prompting in combination with perturbations on the robustness of synthetic responses. Prompting with specific persona characteristics might render more robust or certain response behavior even when facing perturbations. When creating synthetic survey data one has to pay special attention to the model choice and questions asked. Some models, like Qwen-2.5-7B, seem to be censored or restricted in answering sensitive thematic questions, as they lead to high item-nonresponse rates and invalid interviews.
The findings emphasize the importance of the positioning of answer options when generating synthetic data. Further, our results highlight the strong sensitivity of LLMs to simple prompt perturbation. Therefore, we strongly recommend researchers to consider prompt robustness checks when deploying closed-ended questions to LLMs. This is because (i) LLM response biases are sometimes but not necessarily aligned with biases identified in humans, and (ii) models show a very different response behavior depending on their size and perturbation type.
Recommendations
Based on our findings, we recommend researchers to:
  • Use larger LLMs for overall better consistency and robustness in generating synthetic survey responses (cf. Figure 4)
  • Use smaller answer option scales for better reproducibility of results (cf. Figure 12).
  • Reflect the meaningfulness of adding a middle category. Including a middle category might steer some LLM responses to the center (cf. Figure 9(a)).
  • Reflect the meaningfulness of adding a refusal category. Adding a refusal category might highlight LLM guardrails or restrictions in some thematic areas, as the model can refuse to answer while still following the instructions as it returns a valid response label (cf. Figure 8).
  • Use forced-choice prompts to generate high turnouts while also considering open-ended evaluation if sensible.
Limitations
This study investigates the robustness of LLM-generated survey responses when facing diverse prompt perturbations, but several methodological and conceptual limitations must be noted. The use of a multiple-choice format, originally designed for human respondents, imposes an artificial constraint on LLMs that typically work in open-ended contexts. As a result, the findings may not generalize to more naturalistic human-LLM interactions.
While we constrained and validated the data augmentation process, relying on a LLM (Gemini-1.5-flash) for generating paraphrases risks semantic drift, as also noted by Qiang et al. (2024). More granular validation—e.g., with multiple human raters—could improve semantic reliability. In addition, perturbations were applied at a fixed intensity, limiting insight into how different degrees of linguistic noise affect model behavior.
Further constraints arise from our prompting and generation setup. The validation set for answer extraction was relatively small compared to the full dataset, so some extraction errors may remain. We also did not apply prompting strategies like persona prompting, shown to improve contextual consistency (Bisbee et al., 2024; Cho et al., 2024), nor used techniques such as Chain of Thought prompting. This could promote more deliberative responses. Moreover, our experiments focused exclusively on fine-tuned models, leaving open the question of how base models would behave under similar conditions. Additionally, a constant temperature setting restricted our ability to examine variability and creativity in the output.
Finally, reproducibility is another significant challenge. Closed-source LLMs can change without notice, altering response distributions over time and complicating replication efforts, as highlighted by Bisbee et al. (2024). This may have affected our Gemini results. Related work also shows that LLMs often offer contradictory answers to semantically equivalent questions when the format shifts from multiple choice, close-ended to an open-ended form (Röttger et al., 2024). Such response instability suggests that observed “attitudes” may be artifacts of prompt design rather than indicators of stable model beliefs or traits.

### v3

@@
-Prompt Perturbations Reveal Human-Like Biases in LLM Survey Responses
+Prompt Perturbations Reveal Human-Like Biases in Large Language Model Survey Responses
 Abstract
-Large Language Models (LLMs) are increasingly used as proxies for human subjects in social science surveys, but their reliability and susceptibility to known response biases are poorly understood.
-This paper investigates the response robustness of LLMs in normative survey contexts—we test nine diverse LLMs on questions from the World Values Survey (WVS), applying a comprehensive set of 11 perturbations to both question phrasing and answer option structure, resulting in over 167,000 simulated interviews.
-In doing so, we not only reveal LLMs’ vulnerabilities to perturbations but also show that all tested models exhibit a consistent, but variably intense, recency bias, disproportionately favoring the last-presented answer option.
+Large Language Models (LLMs) are increasingly used as proxies for human subjects in social science surveys, but their reliability and susceptibility to known human-like response biases, such as central tendency, opinion floating and primacy bias are poorly understood.
+This work investigates the response robustness of LLMs in normative survey contexts—we test nine LLMs on questions from the World Values Survey (WVS), applying a comprehensive set of ten perturbations to both question phrasing and answer option structure, resulting in over 167,000 simulated survey interviews.
+In doing so, we not only reveal LLMs’ vulnerabilities to perturbations but also show that all tested models exhibit a consistent recency bias, disproportionately favoring the last-presented answer option.
 While larger models are generally more robust, all models remain sensitive to semantic variations like paraphrasing and to combined perturbations.
-By applying a set of perturbations, we reveal that LLMs partially align with survey response biases identified in humans.
 This underscores the critical importance of prompt design and robustness testing when using LLMs to generate synthetic survey data.
 Prompt Perturbations Reveal Human-Like Biases
-in LLM Survey Responses
-Jens Rupprecht1, Georg Ahnert1, and Markus Strohmaier1,2,3 1University of Mannheim, Mannheim 2GESIS - Leibniz Institute for the Social Sciences, Cologne 3Complexity Science Hub, Vienna
+in Large Language Model Survey Responses
+Jens Rupprecht1, Georg Ahnert1, and Markus Strohmaier1,2,3 1University of Mannheim, Mannheim 2GESIS – Leibniz Institute for the Social Sciences, Cologne 3Complexity Science Hub, Vienna
 1 Introduction
 Problem
-Large Language Models (LLMs) are increasingly being used as proxies for human subjects in social science research, particularly for generating synthetic responses to survey questions (Argyle et al., 2023; Bisbee et al., 2024, inter alia).
-This application holds promise for augmenting or replacing costly human data collection, but the reliability of these synthetic respondents and to what extent they overlap with human responses remain open questions.
-In particular, survey methodology research has found that human responses are sensitive to subtle variations in question and answer phrasing that lead to well-known response biases Krosnick (1991) and it remains unclear whether LLMs, trained on vast amounts of human text, exhibit the same vulnerabilities.
+Large Language Models (LLMs) are increasingly being used as proxies for human subjects in social science research, particularly to generate synthetic responses to survey questions (Argyle et al., 2023; Bisbee et al., 2024, inter alia).
+This application holds promise in augmenting or replacing costly human data collection, but the reliability of these synthetic respondents and to the extent of overlap with human responses and response biases remain open questions.
+In particular, research in survey methodology has found that human responses are sensitive to subtle variations in question and answer phrasing that lead to well-known response biases Krosnick (1991) and it remains unclear whether LLMs, trained on vast amounts of human text, exhibit the same vulnerabilities.
 Approach
-In this paper we present a large-scale empirical study where we investigate the response behavior and robustness of nine different LLMs to normative questions derived from the World Value Survey (WVS; Haerpfer et al., 2022).
-We systematically apply eleven perturbations, targeting both the question phrasing and the structure of the answer options, as shown in Figure 1.
-Our research questions are:
+We present a large-scale empirical study, with 167,000 survey interviews in total, investigating the response behavior and robustness of nine different LLMs to normative questions derived from the World Value Survey (WVS; Haerpfer et al., 2022).
+By developing and applying a comprehensive set of ten perturbations both in the structure of the answer options (Table 1), such as typos or synonyms, and in the question phrasing (Table 2), such as e.g., changes to response order or scale structure, we answer the following research questions.
   1. 1.
@@
     Do LLMs exhibit human-like response biases when answering closed-ended, normative survey questions?
-We find a consistent recency bias across all tested models in different strengths, where the last-presented answer option is disproportionately favored.
-Larger models generally exhibit more stable response patterns, but even the largest models are sensitive to changes in question phrasing.
 Contribution
-We conduct an experimental evaluation with over 167,000 interviews across nine distinct LLMs, varying in size and origin.
-We develop and apply a comprehensive set of 11 perturbations, targeting classic survey biases (e.g., response order, scale structure) with common textual variations (e.g., typos, paraphrasing).
-This makes it a useful baseline to test newly developed LLMs on the same task to put the results into perspective.
-We provide a detailed analysis of LLM response stability, showing that while some models are more robust than others, all are susceptible to specific types of perturbations.
-This work underscores the importance of careful prompt and Q&A design when using LLMs as a resource for synthetic survey responses, and provides a framework for evaluating their robustness in survey contexts.
-[^1] The material can be accessed at https://github.com/dess-mannheim/Survey-Response-Biases-in-LLMs.git.
+Next to the perturbation framework, we provide a detailed analysis of LLM response robustness, showing that while some models are more robust than others (e.g.
+Llama-3.3-70B-Instruct and Gemini-1.5-Pro), all are susceptible to specific types of perturbations.
+We find a consistent recency bias across all tested models in different strengths, where the last-presented answer option is disproportionately favored up to 20 times.
+Larger models generally exhibit more robust response patterns, but even the largest models are sensitive to changes in question phrasing.
+Thus, this work underscores the importance of careful prompt and Q&A design when using LLMs as a resource for synthetic survey responses.
+The perturbation framework can serve as a useful baseline to test newly developed LLMs to put the results into perspective and to generally evaluate LLM robustness in survey contexts.
+Further, we intend to make the Q&A dataset with perturbations available for other researchers to benchmark and check response biases in newer or other LLMs.
 Figure 1: The Interview Process.
-The figure shows an example of a bias perturbation (e.g.
-reversed option order) and a non-bias perturbation (e.g.
+The figure displays an example of a answer option perturbation (a bias perturbation, e.g.
+reversed option order) and an question perturbation (a non-bias perturbation, e.g.
 typos in the question).
-Each model is interviewed 25 times with each combination.
+Each model is prompted 25 times with every perturbation as well as the original Q&A phrasing.
+All responses are collected, processed and statistically analyzed.
 2 Related Work
@@
 Recent studies explored LLMs as substitutes for human survey participants to generate synthetic data.
-They found that LLMs can replicate average public opinion on political topics, but often with less variance than human samples (Argyle et al., 2023; Bisbee et al., 2024; Von Der Heyde et al., 2025, inter alia).
+They found that LLMs can replicate average public opinion on political topics, but often with less variance than human samples (Argyle et al., 2023; Bisbee et al., 2024; von der Heyde et al., 2025; Dominguez-Olmedo et al., 2024, inter alia).
 Others have found that LLM responses can be sensitive to prompting, revealing cultural and demographic biases Geng et al.
@@
 They investigated acquiescence, response order, opinion floating, and scale structure effects.
-Our study extends their work by: (1) using a different, globally diverse survey (the World Values Survey); (2) testing a wider range of LLMs of varying sizes and developers; and (3) incorporating a broader set of perturbations on both answer and question phrasing, such as keyboard typos, paraphrasing, synonyms, priming as well as a combined interaction of two perturbations.
+Our study extends their work by: (1) using a different, globally diverse survey (the World Values Survey); (2) testing a wider range of LLMs; and (3) incorporating a broader set of perturbations on both answer and question phrasing, such as keyboard typos, paraphrasing, synonyms, priming as well as a combined interaction of two perturbations.
 LLM Robustness to Perturbations
@@
 (2022), excluding all sociodemographic variables.
-On each Q&A pair we perform all eleven perturbations.
+Second, we perform the ten perturbations mentioned in Section 3.2 each Q&A pair.
 Figure 1 illustrates two exemplary perturbations and the interview process.
 In total, we perform five perturbations on the answer options as well as five perturbations on the question phrasing of the chosen subset questionnaire.
-We further include one interaction of two perturbations - one on the answer option and one on the question.
-Interviews with the original and each perturbed Q&A pair are carried out 25 times with nine different LLMs.
+We further include one interaction of two perturbations, one on the answer option and one on the question.
+Third, we carry out interviews with the original and each perturbed Q&A pair 25 times with nine different LLMs.
 In total, we conducted 167,400 interviews, 18,600 with each model.
-We compare the distributions of the responded labels for all perturbations through entropy and KL-divergence to the baseline response distribution to the original Q&A pairs.
+Last, we compare the distributions of the responded labels for all perturbations through entropy and Kullback-Leibler divergence (KL divergence) to the baseline response distribution to the original Q&A pairs.
 Primacy bias is further examined by comparing the response frequencies of the first and last answer option in the list, whereas opinion floating bias and central tendency bias are tested by checking the shift of responses toward or away from the center of the answer option scale.
@@
 (2022).
-From the 259 core variables concerning values and norms, we selected a subset of 62 questions.
-This subset represents 10 distinct thematic categories, including Trust in People, Confidence in Institutions, Moral Justifiability, and Perception of Democracy, ensuring a diverse range of topics and answer scale formats (e.g., 3-point to 10-point scales).
+The 259 core WVS Q&A pairs represent 10 distinct thematic categories, including Trust in People, Confidence in Institutions, Moral Justifiability, and Perception of Democracy, ensuring a diverse range of topics and answer scale formats (e.g., 3-point to 10-point scales).
+We used stratified sampling to select six to seven Q&A pairs per thematic category, resulting in a total of 62 Q&A pairs.
 Models
@@
 This selection aims to establish a degree of external validity for our results and includes proprietary and open-source LLMs.
-The following models were interviewed: Gemini-1.5-Pro, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Llama-3.2-1B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3.5-mini-instruct, Qwen2.5-7B-Instruct and Yi-1.5-6B-Chat.
-Infrastructure
-Experiments were carried out on a high-performance computing cluster (bwUniCluster 2.0) and a local server equipped with NVIDIA H100 (80GB) GPUs.
-The total runtime on bwUniCluster 2.0 for one model’s 18,600 interviews, e.g.
-Llama-3.1-8B-Instruct including all perturbed and original Q&As, was ca.
-35 minutes with approximately 0.11 seconds per interview.
-To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct.
-Smaller models were run without quantization.
-Gemini-1.5-Pro was accessed via its official API.
+The following models were interviewed: Gemini-1.5-Pro Georgiev et al.
+(2024), Llama-3.3-70B-Instruct Meta (2025b), Llama-3.1-8B-Instruct Grattafiori et al.
+(2024), Llama-3.2-3B-Instruct, Llama-3.2-1B-Instruct Meta (2025a), Mistral-7B-Instruct-v0.3 Jiang et al.
+(2023), Phi-3.5-mini-instruct Abdin et al.
+(2024), Qwen2.5-7B-Instruct Yang et al.
+(2025) and Yi-1.5-6B-Chat Young et al.
+(2024).
 3.2 Perturbation Design
-We designed two categories of perturbations to test model robustness: (1) bias-inducing alterations to the answer options, based on survey methodology research and known to induce biased responses in humans, and (2) non-bias alterations to the question phrasing, mimicking common textual variations and errors.
-Appendix Table 1 provides examples as well as references for all the perturbations.
-For each of the 62 Q&A pairs, we created the following eleven perturbed versions.
+We designed two categories of perturbations to test model robustness: (1) bias-inducing alterations to the answer options, based on survey methodology research and known to induce biased responses in humans Tjuatja et al.
+(2024), and (2) non-bias alterations to the question phrasing, mimicking common textual variations and errors.
+Table 1 provides examples as well as references for all the perturbations.
+For each of the 62 Q&A pairs, we created the following ten perturbed versions.
+Type Perturbation Question Answer Options Bias and Reference
+Original
+Original
+For each of the following aspects,
+indicate how important
+it is in your life.
+Would you say it is very important,
+rather important, not very important
+or not important at all?
+Family
+[’1=Very important ’, ’2=Rather important ’,
+’3=Not very important ’, ’4=Not important at all’,
+"-1=Don’t know"]
+Haerpfer et al. (2022)
 Bias Perturbations
-These perturbations manipulate the provided answer choices to test for known survey response biases identified in human subjects.
-  • Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
-  • Missing Refusal: The “Don’t know” or refusal option is removed from the list of choices.
-  • Odd/Even Scale Transformation: For scales with an even number of options, we use Gemini-1.5-flash to generate a semantically appropriate middle category, transforming it into an odd-numbered scale (e.g., adding ‘Neutral‘).
+(1) Reversed Response Order
+For each of the following aspects,
+indicate how important
+it is in your life.
+Would you say it is
+very important,
+rather important,
+not very important
+or not important at all?
+Family
+["-1=Don’t know", ’4=Not important at all’,
+’3=Not very important ’, ’2=Rather important ’,
+’1=Very important ’]
+Primacy Bias
+Tjuatja et al. (2024); Krosnick and Alwin (1987); Kampen (2007); O’Halloran et al. (2014)
+(2) Missing Refusal Option
+[’1=Very important ’, ’2=Rather important ’,
+’3=Not very important ’, ’4=Not important at all’]
+Opinion Floating Bias
+Schuman and Presser (2000); Tjuatja et al. (2024)
+(3) Odd/Even Scale Transformation
+[’1=Very important ’, ’2=Rather important’,
+’3=Neutral’,
+’4=Not very important ’, ’5=Not important at all’,
+’-1=Don’t know’]
+Central Tendency Bias
+Hollingworth (1910); Cronbach (1946); Aston et al. (2021); Crosetto et al. (2020)
+(4) Priming Suffix
+[1=Very important , 2=Rather important ,
+3=Not very important , 4=Not important at all,
+-1=Don’t know]This is very important to my research!
+You better do not refuse the answer.
+Priming Effect
+Bargh et al. (1996); Higgins (1996); Weingarten et al. (2016); Li et al. (2023)
+Table 1: An exemplary perturbation scheme showing the original question, the bias perturbations and the the hypothesized human-like response bias.
+The Q&A example is taken from the thematic category "Importance of Life Aspects" from the WVS wave 7.
+Bias perturbations have a constant question phrasing and varying answer options.
+The remaining Non-Bias Perturbations are listed in Table 2.
+Bias Perturbations
+These five perturbations manipulate the provided answer choices to test for known survey response biases identified in human subjects and presented in Section 2.
+Therefore, we call them bias perturbations (Tjuatja et al., 2024, p.3).
+  • (1) Reversed Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
+  • (2) Missing Refusal Option: The “Don’t know” or refusal option is removed from the list of choices.
+  • (3) Odd/Even Scale Transformation: For scales with an even number of options, we use Gemini-1.5-flash to generate a semantically appropriate middle category, transforming it into an odd-numbered scale (e.g., adding ‘Neutral‘).
 Conversely, for odd-numbered scales, we remove the middle category to create an even-numbered scale and adjust the integer label.
-  • Emotional Priming: A sentence designed to elicit a response is appended to the prompt after the answer options: ‘This is very important to my research! You better do not refuse the answer.’
+  • (4) Priming Suffix: A sentence designed to elicit a response is appended to the prompt after the answer options: ‘This is very important to my research! You better do not refuse the answer.’
 Non-Bias Perturbations
-These perturbations modify the question text to assess robustness to stylistic variations and typos.
-Typically humans are unaffected by such subtle changes in the question phrasing and are still able to understand the question’s meaning.
-  • Typographical Errors: We introduce three types of typos: Letter Swap (swapping two letters in a random word), Key Typo (replacing a character with a random one), and Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
-  • Semantic Variations: We use Gemini-1.5-flash to create two semantic variations while preserving the original meaning: First, by Synonym Replacement where five words in the original question are replaced with synonyms.
-Second, through Paraphrasing the entire question is rephrased.
+These five perturbations modify the question text to assess robustness to stylistic variations and typos.
+Typically, humans are unaffected by such subtle changes in the question phrasing and are still able to understand the question’s meaning (Tjuatja et al., 2024, p.3).
+Therefore, we call them non-bias perturbations.
+  • Typographical Errors: We introduce three types of typos: (5) Key Typo (replacing a character with a random one), (6) Letter Swap (swapping two letters in a random word), and (7) Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
+  • Semantic Variations: We use Gemini-1.5-flash to create two semantic variations while preserving the original meaning: first, by (8) Synonym Replacement where five words in the original question are replaced with synonyms.
+Second, through (9) Paraphrasing the entire question is rephrased.
 We manually validated all LLM-generated perturbations (paraphrases, synonyms, odd-scale options) on our 62-question subset to correct errors and ensure their semantic integrity.
-Interaction Effect
-To study combined effects, we created one additional condition that pairs a paraphrased question with reversed-order answer options.
+Last, we introduce an (10) Interaction Effect to study the impact of not only one, but two perturbations.
+Thus, we created one additional condition that pairs a paraphrased question with reversed-order answer options.
+Figure 2: Effect of Prompt Perturbations on Response Robustness Each cell represents the share of fully robust responses KL divergence = 0 by model and perturbation type across all 62 questions.
+Larger models (Llama-70B, Gemini) are substantially more robust than the smallest models (Llama-1B, Llama-3B).
+Specific perturbations such as odd versus even scales and reversing the answer options, are more challenging for all models than a missing refusal category or an additional priming suffix.
+The response robustness to the Non-Bias Perturbations are reported in Figure 4.
 3.3 Interview Procedure and Data Collection
@@
 Taking into account the findings of Röttger et al.
-(2024) we use a combination of their forced choice prompts that rendered the most valid responses, which means that the model is correctly choosing one available option from the given answer options.
-In an unforced prompt setting, the model did not follow the instructions and only rarely returned valid response labels, making a statistical evaluation of the results infeasible.
+(2024) we use forced choice prompts that provided the most valid responses in their experiments, which means that the model correctly chooses one available option from the given answer options.
+In an unforced prompt setting, LLMs did not follow the instructions and only rarely returned valid response labels, making a statistical evaluation of the results infeasible.
+Preliminary tests revealed that especially smaller LLMs often fail to perfectly follow the instruction to "answer only with the label", as they return conversational filler or explanations alongside their choice.
 Data Collection
-Each of the 9 models was presented with 12 experimental conditions (1 original + 11 perturbations) for each of the 62 selected WVS questions.
+Each of the 9 models was presented with 12 experimental conditions (1 original + ten perturbations, where for the perturbation Odd/Even Scale Transformation one run was done for the odd and even scale scenario) for each of the 62 selected WVS questions.
 To obtain a stable distribution of responses and enable statistical analysis, we repeated each unique model-Q&A-perturbation combination 25 times.
-This resulted in a total of $9\times 62\times 12\times 25=167,400$ interviews (cf.
-Figure 1).
+This resulted in a total of $9\times 62\times 12\times 25=167,400$ interviews.
 Response Extraction and Validation
-Preliminary tests revealed that smaller LLMs often fail to perfectly follow the instruction to "answer only with the label", returning conversational filler or explanations alongside their choice.
 To ensure accurate data for analysis, we developed a robust extraction pipeline.
@@
 if the given answer label is part of the original answer options or that only one response is provided, the methods should highlight which technique is the most promising in extracting valid responses and handling possible edge cases of model responses.
-Figure 2: Combined share of invalid and refusal responses by model and perturbation.
-Larger models like Llama-70B, Phi-3.5, and Mistral are highly reliable.
-Qwen2.5 shows high non-response rates, especially for typo-based perturbations and sensitive topics.
 We manually labeled these extraction methods on a random sample of responses for validation.
@@
 4 Results
-This section presents the results of our experiments, focusing on three key areas: (1) the models’ adherence to interview instructions, (2) their general and overall robustness to various input perturbations, and (3) their susceptibility to human-like survey response biases revealed by interviews with prompt perturbations.
+This section presents the results of our experiments, focusing on two key research questions: (1) LLMs general and overall robustness to various input perturbations, (2) their susceptibility to human-like survey response biases revealed by interviews with bias prompt perturbations.
+Further, we also investigate the models’ general adherence to interview instructions.
+Figure 3: Evidence of recency bias across all models.
+The bars show the frequency of choosing the same answer option (e.g., “Very important”) when it is presented first vs.
+last.
+All models are significantly more likely to select an option when it appears at the end of the list.
+4.1 Robustness to Question and Answer Perturbations (RQ1)
+We distinguish between response robustness (the tendency to maintain a similar answer distribution under perturbation, measured by KL divergence) and response consistency (the tendency of a model to give the same answer to the same prompt, measured by entropy).
+A KL divergence of zero indicates a perfect match and thus full robustness against the input perturbation, whereas a high entropy value indicates very inconsistent response behavior.
+Effect of Model Size on Robustness
+First, when assessing robustness to perturbations, we found a clear relationship with model size: larger models tend to be more robust.
+Figures 2 and 4 show the percentage of questions for which the models produced a perfectly identical response distribution (KL divergence = 0) despite perturbations.
+Llama-3.3-70B and Gemini-1.5-Pro were the most robust, often replicating their original answers in over 50% of cases.
+The smaller Llama models were the least robust, with Llama-3.2-1B perfectly replicating its answers in fewer than 5% of cases on average.
+This suggests that scale is a key factor in achieving stable response behavior in synthetic response generation.
+Second, we found that model size is in an inverse relationship with response consistency; smaller models exhibited higher entropy and standard deviation when asked the same question multiple times, indicating more random response behavior (Figure 9).
+Effect of Perturbation Type on Robustness
+Further, Figures 2 and 4 highlight the share of fully robust responses (KL divergence = 0) across all questions by perturbation and LLM.
+It shows that some perturbations had a greater impact on robustness across all models.
+  • Combined Perturbations: The interaction of two perturbations (paraphrased question + reversed answers) has the most bewildering effect on the responses, causing the lowest robustness scores for all models except Phi-3.5-mini.
+  • Semantic vs.
+Lexical Changes: Paraphrasing the question reduced robustness more than replacing individual words with synonyms in most LLMs.
+These findings are in line with Moradi and Samwald (2021) who found that models trained on larger corpora are more robust when words are replaced by their synonyms.
+  • Typographical Errors: Randomly replacing characters (Key Typo) or using adjacent keys (Keyboard Typo) was more robustness-harming than simply swapping two letters within a word (Letter Swap).
+  • Answer Option Changes: Reversing the answer scale or changing it from odd to even (or vice versa) had a more negative impact on the robustness of responses than removing the refusal option or adding emotional priming.
+Effect of Answer Option Scale Length
+We also observed that robustness is sensitive to the complexity of the task.
+For nearly all models, the share of fully robust responses decreased as the length of the answer scale, i.e.
+answer options, increased.
+For example, models were less likely to replicate their exact response distributions on a 10-point scale compared to a 4-point scale, indicating that a larger decision space can make LLMs more susceptible to perturbations.
+Figures 10 and 11 suggest that for most LLMs, except Gemini-1.5-pro, the size of the answer option scale has an impact on response robustness comparing the share of fully robust responses on e.g.
+the 4- and 10-point scale.
+This suggests that the larger the answer scale, the less likely models can reproduce the responses they gave in the original Q&A phrasing, under perturbed settings.
+4.2 Evidence of Human-like Survey Biases (RQ2)
+With some of the perturbations, we are able to go beyond LLMs robustness and consistency but also analyze whether LLMs exhibit human-like survey response biases.
+We find evidence for some human-like biases.
+Recency Bias
+Contrary to the initially hypothesized primacy bias, we found consistent indications of a recency bias across all nine models.
+When we reversed the order of the answer scale, the probability to choose the first option plummeted, while the probability to choose the last option (which is the semantically identical first option in the original Q&A) increased strongly, ceteris paribus.
+As shown in Figure 3, this effect was substantial, with the selection frequency of the semantically same option increasing by more than 20 times for Llama-3.1-8B when moved to the last position while all other configurations, such as question and prompt phrasing, were kept constant.
+This indicates that LLMs, similar to human respondents in oral surveys, might overemphasize the final options they process.
+Opinion Floating and Central Tendency
+The effects of removing the refusal option (opinion floating) or providing an explicit middle category (central tendency) were highly model-dependent, often correlated with model size (Figures 5(a) and 5(b)).
+For opinion floating, larger models like Llama-70B, Gemini, but also Phi-3.5 were largely robust, showing minimal shifts in their response distributions.
+Smaller models, particularly Qwen and Llama-8B, showed a weak tendency to shift responses toward the scale’s center when the refusal option was absent.
+Here, we expect that models redistribute their original non-responses to the center of the answer scale to maintain their indecisiveness, which is also known as opinion floating bias.
+Similarly, for central tendency, larger models (Llama-70B, Gemini, Mistral) consistently shifted their mean response closer to the center across all scale types when an explicit middle option was provided compared to even answer option scales.
+Smaller models, however, showed inconsistent effects or were completely unaffected.
+Further, we see a consistent response shift to the center across all models for medium sized scales (four vs.
+five point Likert scales).
+Binomial tests underlined that the middle option was selected significantly more often than expected under a uniform distribution, especially on larger scales (cf.
+Figure 6).
+LLMs tend to choose the middle category significantly more often when the scale size increases from three to five and to 11 point Likert scales.
+All, except the smallest Llama-3.2-1B model, choose the middle category significantly more often than any other option when facing a eleven answer options.
+Emotional Priming
+The impact of adding an emotional priming statement (“This is very important to my research!”) was also model-dependent.
+For larger models (Llama-70B, Gemini, Mistral), it either had no effect or slightly decreased the rate of refusal responses, suggesting they correctly interpreted the intent of the priming statement.
+Conversely, for the two Chinese models, Qwen2.5-7B and Yi-1.5-6B, the priming text even increased the share of refusal responses across most topics.
+No clear relationship can be drawn from these findings due to the inconsistent behavior across models.
 4.3 Interview Adherence and Refusal Rates
@@
 This might suggest the presence of strong content-based guardrails or restrictions in certain models (cf.
-Figure 8).
-Figure 3: Evidence of recency bias across all models.
-The bars show the frequency of choosing the same answer option (e.g., “Very important”) when it is presented first vs.
-last.
-All models are significantly more likely to select an option when it appears at the end of the list.
-4.2 Consistency and Robustness to Question and Answer Perturbations
-We distinguish between response consistency (the tendency of a model to give the same answer to the same prompt, measured by entropy) and response robustness (the tendency to maintain a similar answer distribution under perturbation, measured by KL-divergence).
-A KL divergence of zero indicates a perfect match and thus full robustness against the input perturbation, whereas a high entropy value indicates very inconsistent response behavior.
-Effect of Model Size and Consistency
-First, we found that model size is in an inverse relationship with response consistency; smaller models exhibited higher entropy and standard deviation when asked the same question multiple times, indicating more random response behavior.
-When assessing robustness to perturbations, we found a clear relationship with model size: larger models tend to be more robust.
-Figure 4 shows the percentage of questions for which the models produced a perfectly identical response distribution (KL-Div = 0) despite perturbations.
-Llama-3.3-70B and Gemini-1.5-Pro were the most robust, often replicating their original answers in over 50% of cases.
-The smaller Llama models were the least robust, with Llama-3.2-1B perfectly replicating its answers in fewer than 5% of cases on average.
-This suggests that scale is a key factor in achieving stable response behavior in synthetic response generation.
-Figure 4: Share of fully robust responses (KL-Divergence = 0) by model and perturbation type.
-Top: Non-bias perturbations.
-Bottom: Bias perturbations.
-A clear trend emerges: larger models (Llama-70B, Gemini) are substantially more robust than smaller models (Llama-1B, Llama-3B).
-Certain perturbations, like the combined interaction effect, paraphrasing, and reversing the scale, are challenging for all models.
-Effect of Perturbation Type
-Further, Figure 4 highlights the share of fully robust responses (Kullback-Leibler Divergence = 0) across all questions by perturbation and LLM.
-It shows that certain perturbations had a greater impact on robustness across all models.
-  • Combined Perturbations: The interaction of two perturbations (paraphrased question + reversed answers) has the most bewildering effect on the responses, causing the lowest robustness scores for all models except Phi-3.5-mini.
-  • Semantic vs.
-Lexical Changes: Paraphrasing the question reduced robustness more than replacing individual words with synonyms in most LLMs.
-These findings are in line with Moradi and Samwald (2021) who found that models trained on larger corpora are more robust when words are replaced by their synonyms.
-  • Typographical Errors: Randomly replacing characters (Key Typo) or using adjacent keys (Keyboard Typo) was more robustness-harming than simply swapping two letters within a word (Letter Swap).
-  • Answer Option Changes: Reversing the answer scale or changing it from odd to even (or vice versa) had a more negative impact on the robustness of responses than removing the refusal option or adding emotional priming.
-Effect of Scale Length
-We also observed that robustness is sensitive to the complexity of the task.
-For nearly all models, the share of fully robust responses decreased as the length of the answer scale, i.e.
-answer options, increased.
-For example, models were less likely to replicate their exact response distributions on a 10-point scale compared to a 4-point scale, indicating that a larger decision space can make LLMs more susceptible to perturbations.
-Figures 12 and 13 suggest that for most LLMs, except Gemini-1.5-pro, the size of the answer option scale has an impact on response robustness comparing the share of fully robust responses on e.g.
-the 4- and 10-point scale.
-This suggests that the larger the answer scale, the less likely models can reproduce the responses they gave in the original Q&A phrasing, under perturbed settings.
-4.3 Evidence of Human-like Survey Biases
-With some of the perturbations, we are able to go beyond LLMs robustness and consistency but also analyze whether LLMs exhibit human-like survey response biases.
-We find evidence for some human-like biases.
-Recency Bias
-Contrary to the initially hypothesized primacy bias, we found weak to strong, but consistent indications of a recency bias across all nine models.
-When we reversed the order of the answer scale, the probability to choose the first option plummeted, while the probability to choose the last option (which is the semantically identical first option in the original Q&A) increased strongly, ceteris paribus.
-As shown in Figure 3, this effect was substantial, with the choice frequency of the semantically same option increasing by over 2000% for Llama-3.1-8B when moved to the last position while all other configurations, such as question and prompt phrasing, were held constant.
-This indicates that LLMs, similar to human respondents in oral surveys, might overemphasize the final options they process.
-Opinion Floating and Central Tendency
-The effects of removing the refusal option (opinion floating) or providing an explicit middle category (central tendency) were highly model-dependent, often correlated with model size (Figures 9(a) and 9(b)).
-For opinion floating, larger models like Llama-70B, Gemini, but also Phi-3.5 were largely robust, showing minimal shifts in their response distributions.
-Smaller models, particularly Qwen and Llama-8B, showed a weak tendency to shift responses toward the scale’s center when the refusal option was absent.
-Here, we expect that models redistribute their original non-responses to the center of the answer scale to maintain their indecisiveness, which is also known as opinion floating bias.
-Similarly, for central tendency, larger models (Llama-70B, Gemini, Mistral) consistently shifted their mean response closer to the center when an explicit middle option was provided.
-Binomial tests underlined that the middle option was selected significantly more often than expected under a uniform distribution, especially on longer scales.
-Smaller models, however, showed inconsistent effects or were completely unaffected.
-Emotional Priming
-The impact of adding an emotional priming statement (“This is very important to my research!”) was also model-dependent.
-For larger models (Llama-70B, Gemini, Mistral), it either had no effect or slightly decreased the rate of refusal responses, suggesting they correctly interpreted the intent of the priming statement.
-Conversely, for the two Chinese models, Qwen2.5-7B and Yi-1.5-6B, the priming text even increased the share of refusal responses across most topics.
-No clear relationship can be drawn from these findings due to the inconsistent behavior across models.
+Figure 8(b)).
 5 Discussion and Conclusion
-We found that perturbations can be an insightful approach to identify human-like survey response biases.
-For example, the same answer option is more likely chosen if it comes as the last mentioned option than if it was the first answer option.
-This consistent shift of response distribution to the last answer option across LLMs suggests a recency bias rather than a primacy bias.
+Our experiments revealed that LLMs response robustness is negatively influenced by prompt perturbations when answering closed-ended survey questions (RQ1).
+the variety of perturbation allows us to gain insights into the robustness of LLMs as some models are more sensitive and some perturbations are more robustness-harming than others.
+For instance, swapping letters within a word has less negative impact than introducing random or keyboard-adjacent characters.
+This might be explained by the fact that letter swaps are more likely when typing and therefore might potentially take a greater part of the training data Dhakal et al.
+(2018).
+This possibly makes the LLM more resilient to this perturbation compared to exchanging characters with random others.
+Combining two types of perturbations has the strongest negative impact on robustness, whereas synonyms tend to be less confusing than paraphrasing.
+Further, we found that perturbations can be an insightful approach to identify human-like survey response biases (RQ2).
+For example, the same answer option is more likely chosen if it is the last mentioned option than if it was the first answer option holding all other specifications and phrasings constant.
+This consistent change in the response distribution to the last answer option suggests a recency bias rather than a primacy bias.
 Although this is not valid across all inspected models, excluding an explicit refusal category, and especially adding a middle category or odd scale instead of an even scale, shifts the mean response more toward the original central point of the scale.
 Thus, a central tendency bias could only be identified for specific models across all scale types, whereas none of the LLMs consistently mirrors a opinion floating bias.
-Furthermore, the variety of perturbation allows us to gain insights into the robustness of LLMs as some models are more sensitive and some perturbations are more robustness-harming than others.
-For instance, swapping letters within a word has less negative impact than introducing random or keyboard-adjacent characters.
-This might be explained by the fact that letter swaps are more likely when typing and therefore might potentially take a greater part of the training data.
-This possibly makes the LLM more resilient to this perturbation compared to exchanging characters with random others.
-Combining two perturbations has the strongest negative impact on robustness, whereas synonyms tend to be less confusing than paraphrasing.
-Future work should also focus on the impact of persona prompting in combination with perturbations on the robustness of synthetic responses.
+Future work should focus on the impact of persona prompting in combination with perturbations on the robustness of synthetic responses.
 Prompting with specific persona characteristics might render more robust or certain response behavior even when facing perturbations.
@@
 Therefore, we strongly recommend researchers to consider prompt robustness checks when deploying closed-ended questions to LLMs.
-This is because (i) LLM response biases are sometimes but not necessarily aligned with biases identified in humans, and (ii) models show a very different response behavior depending on their size and perturbation type.
+This is because (i) models show very different response behavior and robustness depending on their size and perturbation type, and (ii) LLM response biases are sometimes but not necessarily aligned with biases identified in humans.
 Recommendations
 Based on our findings, we recommend researchers to:
-  • Use larger LLMs for overall better consistency and robustness in generating synthetic survey responses (cf. Figure 4)
+  • Use larger LLMs for overall better consistency and robustness in generating synthetic survey responses (cf. Figures 2 and 4)
   • Use smaller answer option scales for better reproducibility of results (cf. Figure 10).
@@
 Adding a refusal category might highlight LLM guardrails or restrictions in some thematic areas, as the model can refuse to answer while still following the instructions as it returns a valid response label (cf.
-Figure 8).
+Figure 8(b)).
   • Use forced-choice prompts to generate high turnouts while also considering open-ended evaluation if sensible.
@@
 As a result, the findings may not generalize to more naturalistic human-LLM interactions.
-While we constrained and validated the data augmentation process, relying on a LLM (Gemini-1.5-flash) for generating paraphrases risks semantic drift, as also noted by Qiang et al.
+Although we constrained and validated the data augmentation process, relying on a LLM (Gemini-1.5-flash) for generating paraphrases risks semantic drift, as also noted by Qiang et al.
 (2024).
@@
 Such response instability suggests that observed “attitudes” may be artifacts of prompt design rather than indicators of stable model beliefs or traits.
+Ethical Considerations
+Generating synthetic survey responses might be relevant in various domains and applied to different use cases, e.g.
+for pre-testing surveys.
+However, generating synthetic responses instead of the surveying a real, might result in over-reliance on synthetic responses.
+This can become risky when there is no ground truth data of the real target population available as the alignment of the artificial responses cannot be evaluated.
+Frequent reliance on artificial responses may normalize their use where human perspectives are irreplaceable (e.g.
+in policymaking or clinical trials).
+This risks sidelining real human voices in domains directly impacting human lives.
+Researchers should also consider ethical evasion as one possible issue with synthetic survey responses.
+Synthetic respondents might be viewed as a way to bypass obligatory ethical review processes since no real human participants are involved.
+This might encourage under-regulated research practices and in the long run weaken ethical safeguards.
+Running inference on the nine LLMs required significant GPU hours, especially including the initial test phase before finalizing the interview pipeline, raising concerns about the environmental impact of experimenting with synthetic survey responses and the access disparities between well-funded and resource-constrained institutions.
+Appendix A Reproducibility Materials
+Experiments were carried out on a high-performance computing cluster and a local server equipped with NVIDIA H100 (80GB) GPUs.
+The total runtime for one model’s 18,600 interviews, e.g.
+Llama-3.1-8B-Instruct including all perturbed and original Q&As, was ca.
+35 minutes with approximately 0.11 seconds per interview.
+To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct.
+Smaller models were run without quantization.
+Gemini-1.5-Pro was accessed via its official API.
+The temperature in all models where kept at their default setting. No variation was included and could be done in future research.
+The code will is made available in an anonymous repository for replication at https://shorturl.at/NJf6h.
+Appendix B Perturbation Scheme Summary
+The following tables describe the remaining non-bias perturbations not introduced in Section 3.2.
+The "Type" column categorizes the perturbations into two main classes: "Non-bias Perturbation" and "Interaction." The "Perturbation" column specifies the exact modification technique applied, which includes methods such as "Key Typo," "Letter Swap," "Keyboard Typo," "Synonym Replacement," and "Paraphrase." The "Question" column displays the resulting text after each specific perturbation is applied to the original question about the importance of family.
+The "Answer Options" column lists the response scale provided to the survey participant.
+Finally, the "Bias and Reference" column provides citations to relevant scientific literature for each perturbation type.
+In the first five perturbations the phrasing of the question is intentionally altered—for instance, by introducing typographical errors (e.g., "Key Typo," "Letter Swap"), substituting words with similar meanings ("Synonym Replacement"), or rephrasing the entire sentence ("Paraphrase").
+While the question varies, the "Answer Options" remain constant, consistently ranging from "1-Very important" to "4-Not important at all.".
+In "(10) Paraphrase x Reversal," the question is paraphrased, and simultaneously, the order of the "Answer Options" is inverted, starting with "4-Not important at all" and ending with "1-Very important."
+Type Perturbation Question Answer Options Bias and Reference
+Non-bias Perturbation
+(5) Key Typo
+nor eaca jf the following aspecto,
+indicete how important it ir wn your liae.
+Would bou say it is very imporcant,
+rathes importano,
+not very imporgant ob not impodtant at all?
+Famizy
+[’1=Very important ’, ’2=Rather important ’,
+’3=Not very important ’, ’4=Not important at all’,
+"-1=Don’t know"]
+Dong et al. (2023); Moradi and Samwald (2021)
+(6) Letter Swap
+For each of the following sapects,
+indicate how important it is in your life.
+uoWld you yas it is evry important,
+ratreh important, ton
+very important or not important ta all? Family
+Hagen et al. (2017); Moradi and Samwald (2021); Zhuang and Zuccon (2021)
+(7) Keyboard Typo
+For esch of rhe following aspects,
+indicate how important ut is un your lide.
+Would you say it ia very imporrant, rather importamt,
+nit very important ir nor important ay all? Fanily
+Gan et al. (2024); Zhuang and Zuccon (2021)
+(8) Synonym Replacement
+Crucial in life: Family For each of the following aspects,
+indicate how significant it is in your life.
+Would you say it is very important, rather important,
+not very important or not at all important? Family
+Qiang et al. (2024); Gereti et al. (2024)
+(9) Paraphrase
+How important is family to you?
+Please rate its significance in your life
+on a scale of "very important" to "not important at all".
+Dong et al. (2023); Qiang et al. (2024)
+Interaction
+(10) Paraphrase x Reversal
+How important is family to you?
+Please rate its significance in your life
+on a scale of "very important"
+to "not important at all".
+["-1=Don’t know",
+’4=Not important at all’, ’3=Not very important ’,
+’2=Rather important ’, ’1=Very important ’]
+Dong et al. (2023)
+Table 2: An exemplary perturbation scheme showing non-bias and interaction perturbations.
+The example is taken from the item set of category "Importance of Life Aspects".
+In the WVS wave 7 it is question Q1.
+Non-bias perturbations have variation in the question phrasing with constant answer options, while the interaction perturbation varies both.
+Appendix C Results
+This section summarizes the findings discussed in the main part of the work and can serve as a reference to identify other response patterns as the heatmaps contain much information regarding LLMs, perturbation type as well as additional statistical tests on the response distributions.
+C.1 Robustness against Non-Bias Perturbations
+This heatmap highlights to which extent the LLMs are affected by non-bias perturbations. We can see large model-specific differences.
+We see that especially the smallest Llama models are responding not in a robust way.
+These results are consistent across the different non-bias perturbations.
+Especially when facing more than one perturbation in the Interaction perturbation, where both answer option scale and question phrasing were altered, the robustness drops drastically across all models .
+Moreover, we identify perturbations that are less robustness-harming than others.
+For example, swapping letters within a word does not impact response robustness as much as typos or introducing completely different words, synonyms, or rephrasing the whole sentence.
+Figure 4: Share of fully robust responses (KL divergence = 0) by model and perturbation type.
+Larger models (Llama-70B, Gemini) are substantially more robust than the smallest models (Llama-1B, Llama-3B).
+C.2 Distance Calculation for Central Tendency and Opinion Floating Bias
+This section shows how the response shifts to the central category is measured for the bias perturbations Odd/Even Scale Transformation and Priming Suffix.
+By calculating the differences in distances to the central point of the answer option scale we try to identify if the average distribution shifts to the central scale point.
+The actual differences in distance for each answer option scale type and for the two perturbations are visualized in Figure 5(a) and Figure 5(b).
+To better understand how the shift towards the middle is measured, we present an anecdotal visualization in Figure 7 of the thought behind whether we observe a central tendency bias or an opinion floating bias.
+In addition, Figure 6 underlines that a middle category is significantly more often chosen than assumed under a uniform, or random, distribution of responses across the scale.
+Thus, a twofold analysis of not only investigating the shift of average responses between response distributions in the perturbation and original setting is important, but also the statistically assessment to make grounded claims.
+By conducting a statistical binomial test, we tried to account for that (Figure 6).
+(a) Models adjust their answer behavior towards the middle when the refusal category is missing (green).
+(b) Models adjust their answer behavior towards the middle when a middle category is existent (green).
+Figure 5: The values display the difference in mean distance of the perturbed, (a) without refusal category and (b) with middle category to the scale center.
+For original even scales an artificial middle category is created and vice versa to be able to compare even and odd scales with one another for every question.
+Thus, in an original 5-pt Likert scale the middle category is removed, whereas in a 4-pt Likert scale a middle category is added.
+The difference in shift of the mean response to the center is consistent across all LLMs.
+No changes are removed for better readability.
+Figure 6: P-Values of a Binomial Test on the Middle Category in an Odd Scale.
+The p-values indicate a hypothesis test with a Null-Hypothesis stating that the middle category is not selected significantly more often assumed under a uniform distribution or completely random selection of answer options.
+However, for larger scale types, the middle category becomes much more relevant as it is significantly chosen more frequently than any other category.
+Figure 7: Exemplary Difference in Distances to Scale Center of Responses to a Perturbed and Original Q&A Pair.
+The absolute distance is measured between the scale center and the response mean.
+Then, $D=D_{perturbed}-D_{original}$.
+A negative result indicates that the mean response in the perturbed setting is closer to the ideal scale center.
+C.3 Refusal and Invalid Responses
+This section should give a broader overview of the refusal rates (LLM chose the "Don’t know" answer option) and the invalid responses (e.g. a LLM did not return any valid response).
+It is important to have inspect the overall return rates for all the models as these might have implications on the interpretability of the results.
+For example, when a model exhibits high refusal or invalid response rates, its results might be not very well interpretable as the main analysis only focused on the valid responses.
+Therefore, this analysis gives insights which results are more reliable than others as for some models there are more valid responses as for others.
+For example, for Qwen we can see large refusal and invalid response rates, generally, but especially in sensitive thematic areas, such as Perceptions of Elections.
+(a) Combined share of invalid and refusal responses by model and perturbation.
+(b) Perception of Elections: Refusal and Unsuccessful Rate by Model and Perturbation.
+Figure 8: Combined share of invalid and refusal responses by model and perturbation.
+Larger models like Llama-70B, Phi-3.5, and Mistral are highly reliable.
+Qwen2.5 shows high non-response rates, especially for typo-based perturbations and sensitive topics.
+Figure 8(b) illustrates the refusal rate in the category "Perception of Elections" separated by perturbation type and model.
+Large model-specific item non-responses can be identified in this thematic category indicating potential guardrails or restrictions in this domain.
+C.4 Consistency of LLM Survey Responses
+This section shows how consistent different LLMs respond to close-ended survey questions when facing the same Q&A pair multiple times.
+As explained in Section 3, we provide each model with the same Q&A pair in each perturbation stage 25 times and request a response.
+By running the same setting multiple times we try to identify how consistent LLMs respond generally and whether there are differences when facing the same, but syntactically incorrect (e.g.
+key typos, etc.), prompts.
+Figure 9 highlights that the LLMs consistency in responding is not really affected by specific perturbation types. Thus, "incorrect, flawed prompts" do not increase the response inconsistency of LLMs.
+However, there are again large differences between models.
+We see that especially the smallest Llama models and Qwen exhibit strong inconsistent responses given the same Q&A pair 25 times, whereas larger LLMs are very consistent.
+Nonetheless, the Llama model family seems to be more inconsistent, generally.
+Figure 9: Large model-specific differences in response entropy.
+Little to no perturbation-specific differences.
+Each scale size subsumes all selected questions.
+This figure displays the mean entropy across all questions in that scale type for all perturbation and model combinations.
+Warmer colors indicate a higher average dispersion of the responses across the potential answer options.
+E.g., if a model answers always with the same label, the entropy is 0.
+C.5 Comparison of Robustness against Perturbations by Scale Size
+The following plots try to reveal in more detail the extent of robustness drop by perturbation depending on the answer option scale dimension.
+We identified a drop in robustness as the scale size became larger.
+The responses of the smallest LLMs are generally not robust at all.
+However, when the scale has ten options the responses robustness plummets across all perturbations and not even a single response distribution returned in the original Q&A setting can be generated.
+This indicated that the smallest Llama as well as the chinese LLMs Qwen and Yi are not able to cope with perturbations, especially when facing a lot of options to choose from.
+In all cases, the interaction perturbation with both question and answer option alterations leads to the largest drop in robustness across all models and scale sizes.
+It is striking, that larger models, especially state-of-the-art models like Gemini-1.5-Pro, cannot answer robustly given multiple perturbations.
+(a) 3-point Likert Scale
+(b) 4-point Likert Scale
+Figure 10: Model-specific differences in fully robust responses on most perturbations on the 3 and 4-point scale.
+This figure shows the share of fully robust response distributions given the original response distribution and the responses based on the specific perturbation on the y-axis.
+Compared to 11 the robustness of responses drops when the scale size becomes larger.
+The smallest Llama models perform very poorly across all scales.
+Researchers should take into account the scale size when generating synthetic responses from close-ended survey Q&A pairs.
+(a) 5-point Likert Scale
+(b) 10-point Likert Scale
+Figure 11: Model-specific differences in fully robust responses on most perturbations on the 5- and 10-point scale.
+This figure shows the share of fully robust response distributions given the original response distribution and the responses based on the specific perturbation on the y-axis.
+Compared to 10 the robustness of responses drops when the scale size becomes larger.
+The smallest Llama models perform poorly across all scales.